cmss10 scaled \magstep0 22:17 - 4/20/2000

cmss10

0 1 2 3 4 5 6 7

’000

Γ Δ Θ Λ Ξ Π Σ Υ

’010

Φ Ψ Ω ff fi fl ffi ffl
’020 ı ȷ `  ´  ˇ  ˘  ¯  ˚
’030 ¸ ß æ œ ø Æ Œ Ø
’040 ! ” # $ % & ’
’050 ( ) * + , - . /
’060 0 1 2 3 4 5 6 7
’070 8 9 : ; ¡ = ¿ ?
’100 @ A B C D E F G

’110

H I J K L M N O

’120

P Q R S T U V W

’130 X Y Z [ “ ] ˆ  ˙
’140 ‘ a b c d e f g

’150

h i j k l m n o ’160

p q r s t u v w

’170 x y z – — ˝  ˜  ¨

abcdefghijklmnopqrstuvwxyz Alphabet length 119.62526pt. Quad is 10.00002pt Lower case height is 6.94444pt; depth is 1.94444pt

ABCDEFGHIJKLMNOPQRSTUVWXYZ Alphabet length 170.13945pt; height 6.94444pt

The design of low-resolution typefaces is complicated by the need to achieve two different and rather incompatible ends. A font which is exclusively intended for Office Automation equipment should be designed to make the very best use of that environment. Its proportions, balance, shading and letter-spacing can and indeed must be fitted to exact pixel-boundaries, whether or not the refinements of anti-aliasing are used.

A low-resolution font which was created primarily as the proof-copy imitation of an associated high-resolution font will never be quite so precisely matched to its environment. Its contours will necessarily be determined by exact pixel-boundaries, but often not in the way we might prefer. It need not be sloppy, but it can never be so well tuned as a font of the first variety. It is a poor Platonic reflection of an idea that cannot be fully known except in the dazzling light of high-resolution typesetting. The vast majority of fonts produced with METAFONT belong to this second variety. A low-resolution METAFONT character should be judged for what it is, and not for what it has never claimed to be.